The sound of heavy automobile traffic, heard at a certain distance, has a sound level of about 75 dB. What is the intensity of the noise there? answer in W/m^2.
1
Expert's answer
2011-06-22T10:47:20-0400
The measure of a ratio of two sound intensities is
L1=10 log10(I1/I0)
where I1 and I0 are the intensities.
So, we have
75=10 log10(I1/I0 ) → log10(I1/I0 ) = 75/10 → I1/I0 = 107.5 = √(1015 ) W/m2
The sound intensity level is given the letter "L1" and is measured in "dB". The decibel is a dimensionless quantity.